OpenStudy (anonymous):

Find the slope of R= -1+sin theta at the points theta= 0 and theta= pie

OpenStudy (anonymous):

yes, polar

OpenStudy (anonymous):

oops\[\frac{dy}{dx}=\frac{r'\sin(\theta)+r\cos(\theta)}{r'\cos(\theta)-r\sin(\theta)}\]

OpenStudy (anonymous):

in your case \(r=-1+\sin(\theta)\) and \(r'=\cos(\theta)\)

OpenStudy (anonymous):

i got the derivative part, it is just the limits are confusing me, what am i suppose to do with me?

OpenStudy (primeralph):

Plug them in.

OpenStudy (primeralph):

Those are not limits; they are distinct points.

OpenStudy (anonymous):

what @primeralph said, evaluate unless you get \(\frac{0}{0}\)

OpenStudy (anonymous):

oh, so plug both of them into the tangent equation and get two separate tangents??

OpenStudy (primeralph):

Yes.

OpenStudy (primeralph):

Tangents are at points; integrals are over bounds (limits).

OpenStudy (anonymous):

thanks!!

OpenStudy (primeralph):

Well, you're easy to teach. You seem smart already, so keep it up.